9.A certain aircraft has a liftoff speed of 129 km/h.(a) What minimum constant acceleration does the aircraft require if it is to be airborne after a takeoff run of 234 m?

 m/s2


(b) How long does it take the aircraft to become airborne?

 s


1
Expert's answer
2022-06-03T15:25:20-0400

Given:

"v=129\:\rm km/h=35.8\: m/s"

"d=234\:\rm m"

(a)

"a=\frac{v^2}{2d}=\frac{35.8^2}{2*234}=2.74\:\rm m/s^2"

(b)

"t=\frac{v}{a}=\frac{35.8}{2.74}=13.0\:\rm s"